OpenStudy (anonymous):

If Sue uses a force of 80 N to move a box that had a mass of 35 kg a distance of 25 m, how much work did she do?

OpenStudy (johnweldon1993):

\[\large F = ma\] \[\large F = 35kg \times 9.81 m/s^2 = 343.35N\] Now \[\large W = Fd = 343.35N \times 25m = 8583.75 J\] or \[\large 8.57375 kJ\]

OpenStudy (shamim):

The above solution is wrong!!!

OpenStudy (shamim):

U hv to use force F=80N

OpenStudy (shamim):

W=Fd=80*25=?joule

OpenStudy (anonymous):

yeah I think that's right ^^

OpenStudy (amistre64):

F = 80 cos(a) N ... given that a is the angle between the direction of force and the direction of movement.